OpenStudy (anonymous):

2z-3 is a factor of 2z^3 -z^2 +az -3 find a and all zeroes of the cubic. 2z-3 is a factor of 2z^3 -z^2 +az -3 find a and all zeroes of the cubic. @Mathematics

OpenStudy (anonymous):

if 2z - 3 is a factor then f(3/2) = 0 2(3/2)^3 - (3/2)^2 + 3/2 a - 3 = 0 2 * 27/8 - 9/4 + 3a / 2 - 3 = 0 54 - 18 + 12a - 24 = 0 a = -54 + 18 + 24 / 12 a = -1

OpenStudy (mertsj):

Use synthetic division. Put 3/2 on the shelf. 3/2(a+3) must =3. So a must be -1. FEMA ing factor is therefore 2z^2+2z+2. Use the quadratic formula to find that the remaining zeros are (-1+|-sqrt6)/2

OpenStudy (anonymous):

factorising (2z-3)(z^2 + z + 1) = 0 z = 3/2 z = -1 +- sqrt(1 - 4 *1*1) / 2 = [ -1 + sqrt3 i ] / 2 or [ -1 - sqrt3 i] / 2

OpenStudy (anonymous):

what does that sqrt mean?

OpenStudy (anonymous):

ohh square root??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ha thank you so much!!

OpenStudy (anonymous):

if ound the z^2 + z + 1 by long division - i find i get into trouble with synthetic division

OpenStudy (anonymous):

ok you can further simplify the complex roots to -(1/2) +(1/2)i sqrt3 , -(1/2) - (1/2)i sqrt3

OpenStudy (anonymous):

the sqrt refers to the 3 only

OpenStudy (anonymous):

yapp!! :)